### v1

# Direct Quantized Training of Language Models with Stochastic Rounding

Abstract
Although recent quantized Large Language Models (LLMs), such as BitNet, have paved the way for significant reduction in memory usage during deployment with binary or ternary weights, training these models still demands substantial memory footprints. This is partly because high-precision (i.e., unquantized) weight matrices required for straight-through estimation must be maintained throughout the whole training process. To address this, we explore the potential of directly updating the quantized low-precision weight matrices without relying on the straight-through estimator during backpropagation, thereby saving memory usage during training. Specifically, we employ a stochastic rounding technique to minimize information loss caused by the use of low-bit weights throughout training. Experimental results on our LLaMA-structured models indicate that (1) training with only low-precision weights is feasible even when they are constrained to ternary values, (2) extending the bit width to 8 bits results in only a 5% loss degradation compared to BitNet b1.58 while offering the potential for reduced memory usage during training, and (3) our models can also perform inference using ternary weights, showcasing their flexibility in deployment. [^1] Work in progress. Code is available at https://github.com/KYuuto1006/DQT
Direct Quantized Training of Language Models with Stochastic Rounding
Kaiyan Zhao1, Tsuguchika Tabaru2, Kenichi Kobayashi2, Takumi Honda2, Masafumi Yamazaki2, Yoshimasa Tsuruoka1 1The University of Tokyo, Tokyo, Japan 2Fujitsu Limited, Kawasaki, Japan 1{zhaokaiyan1006, yoshimasa-tsuruoka}@g.ecc.u-tokyo.ac.jp 2{tabaru, kenichi, honda.takumi, m.yamazaki}@fujitsu.com
1 Introduction
Large Language Models (LLMs) have become a promising solution for a wide range of Natural Language Processing (NLP) tasks including machine translation (Xu et al., 2024a; Wu et al., 2024), summarization (Zhang et al., 2024a), reasoning (Wei et al., 2024; Kojima et al., 2024; Miao et al., 2024) and question answering (OpenAI, 2022, 2023). However, their development is challenged by the need for vast datasets and substantial computational resources, especially when the size of current LLMs continually grows larger (Duan et al., 2024).
Quantization, which involves converting high-precision parameter matrices into lower-precision formats, emerges as an effective approach for enabling resource-efficient LLMs (Egashira et al., 2024). Traditional quantization methods can be divided into two categories: Post Training Quantization (PTQ) and Quantization Aware Training (QAT). PTQ reduces the bit precision of weight matrices in an already pre-trained LLM (Banner et al., 2019; Frantar et al., 2023), while QAT incorporates quantization during training, enabling the model to adapt to low-bit precision throughout the learning process (Jacob et al., 2018). Numerous studies have shown that models quantized through QAT generally achieve better performance than those using PTQ because PTQ may lead to a mismatch between the high-precision model’s learned representations and the constraints of low-bit precision (Chen et al., 2024; Liu et al., 2024).
Figure 1: Comparison of the training process for BitNet and our modified one. Upper: The training process for BitNet, where the original high precision weights are updated with the straight-through estimator in backward process. Lower: We directly update the low precision weights with stochastic rounding, eliminating the need to quantize the weight matrices in each training step and keeping weight matrices always ternary.
Recently, a QAT method, BitNet (Wang et al., 2023; Ma et al., 2024), makes the first attempt to quantize full-precision (FP32) transformers into binary or ternary models from scratch, with weight matrices constrained to values of {-1, 0, 1} while maintaining competitive performances with unquantized ones. We illustrate the training process of traditional QAT methods such as BitNet in Figure 1 (a). After computing the loss based on the quantized weights and inputs, the loss is backpropagated to update the original high-precision weights via the straight-through estimator (STE) (Bengio et al., 2013). These weights are then re-quantized in each training step. This iterative process is necessary because quantization itself is not differentiable (Chen et al., 2019), requiring special gradient accumulation methods. As a result, the high-precision weight matrices are always maintained throughout the whole training process, which makes traditional QAT inefficient and takes up a lot of extra memory footprints.
Particularly, when scaling to LLMs, this problem may pose significant challenges to QAT’s practical deployment, as maintaining larger weight matrices in LLMs becomes more resource-intensive. For example, maintaining the weights of a 1B LLM alone would require 4GB of memory in the FP32 format, whereas maintaining ternary weights reduces this to 0.2GB. This requirement for memory footprints limits the accessibility of QAT techniques for researchers and organizations with limited computational resources.
In order to address these challenges, we explore Direct Quantized Training (DQT), a modified QAT approach for language models that maintains only low-precision weights throughout the training process. DQT eliminates the reliance on STE by directly updating the quantized low-precision weights during backpropagation, as shown in Figure 1 (b). Specifically, we use stochastic rounding (Von Neumann and Goldstine, 1947), a rounding technique that probabilistically rounds values to the nearest representable precision based on their distance from those values, thereby preserving the low-precision format of the weight matrices after backpropagation and minimizing the information loss caused by using only low-bit weights. The DQT method keeps all weight matrices fixed at n-bit precision throughout the entire training process. With DQT, the model no longer needs to quantize all high-precision weight matrices at each training step, theoretically accelerating the QAT process and reducing memory usage. More importantly, traditional QAT often introduces extra memory overhead due to the updates on high-precision weights, which limits its practicability in training general LLMs. In contrast, the light-weighted memory dependency of DQT enables quantization’s application even for scenarios where computational resources are constrained.
Experimental results on our LLaMA-structured models trained with DQT demonstrate that (1) models trained with DQT are capable of converging even when weight matrices are constrained to ternary values; (2) with 8-bit DQT, our models can achieve performance levels competitive with BitNet b1.58, showing the feasibility of DQT; and (3) it is also possible for models trained with DQT to utilize only ternary weights during inference while delivering performance comparable to BitNet. We assume that DQT could provide new insights on addressing the computational challenges posed by traditional QAT.
2 Related Works
Quantization for deep neural networks has a history spanning nearly ten years, with researchers initially compressing networks to reduce memory usage and computational load while maintaining accuracy (Rastegari et al., 2016; Hubara et al., 2018). Recent quantization methods for LLMs can be divided into two categories: Post Training Quantization (PTQ) and Quantization Aware Training (QAT).
2.1 Post Training Quantization
PTQ transforms high-precision parameters into low-bit ones after the parameters are already pre-trained. Some approaches utilize a small set of calibration data to accomplish this transformation while preserving the model’s performance (Nagel et al., 2020; Li et al., 2021; Frantar et al., 2023). Others explore methods that eliminate the need for calibration data altogether (Cai et al., 2020). The challenge of PTQ lies in achieving a balance between compression efficiency and minimal accuracy degradation, often involving computational trade-offs that make it a non-trivial task for general-purpose applications. Moreover, the performance of PTQ consistently lags behind that of QAT, as there is a gap between the learned high-precision representations and the constrained bit width (Chen et al., 2024; Liu et al., 2024).
2.2 Quantization Aware Training
To bridge the gap between high-precision parameter training and quantization, QAT incorporates the quantization of model parameters during the training process. However, the substantial computational cost of applying QAT to LLMs has hindered its widespread exploration.
The first application of QAT to LLMs comes from Liu et al. (2024), who propose a data-free distillation-based method and quantize the model to 4 bits. Xu et al. (2024b) further expand distillation-based methods to binary quantization by introducing limited trainable vectors. More recently, BitNet (Wang et al., 2023; Ma et al., 2024) is proposed, achieving training from scratch QAT with weight values constrained to {-1,0,1}.
However, due to the non-differentiability of the quantization process, special gradient approximation methods like the straight-through estimator (STE) (Bengio et al., 2013) are commonly employed during training. While effective, this approach often results in slower training processes and increased computational memory overhead since high-precision weights are always maintained during the training process, limiting its scalability to larger models.
These inefficiencies become particularly pronounced when scaling QAT to larger models, such as those in the billion-parameter range, where the memory and computational demands can quickly surpass the capabilities of standard GPUs, limiting the practical application of QAT for real-world use cases. To address this, in this work we explore the potential for more efficient and practical QAT methods that significantly reduce memory usage compared to traditional QAT approaches.
3 Method
In this section, we first introduce stochastic rounding, the core idea of our proposed method that maintains weight matrices at low-precision during training. Then we move on to describe how stochastic rounding is applied in the modified training process for QAT.
3.1 Stochastic Rounding
The idea of stochastic rounding (SR) originates from Von Neumann and Goldstine (1947), and is initially used for reducing bias in numerical computations and recently for deep learning models (Gupta et al., 2015). It is a rounding technique that probabilistically rounds values to the nearest representable precision based on their distance from those values. Given a high precision value $x$, stochastic rounding can be defined as the following equation Gupta et al. (2015); Zhang et al. (2024b):
$\mathrm{SR}(x)=\left\{\begin{aligned} &\lfloor x\rfloor,\ \mathrm{with}\ p=\lceil x\rceil-x\\ &\lceil x\rceil,\ \mathrm{otherwise}\\ \end{aligned}\right.,$ (1)
where $p$ stands for the probability of turning $x$ to $\mathrm{floor}(x)$ or $\mathrm{ceil}(x)$. In this way, we can naturally quantize high-precision values into low-precision ones.
3.2 Modified Training Process
Next, we continue to explain the details of DQT. As shown in Figure 1 (b), we start training from low-precision weight matrices instead of high-precision ones. We achieve this through utilizing AbsMean Quantization following (Ma et al., 2024) on a randomly initiated weight matrix $W$. The absolute mean value for $W$ can be represented as:
$\mathrm{AbsMean}(W)=\frac{1}{k}\sum_{i=1}^{k}\ \lvert w_{i}\rvert,$ (2)
where $w_{i}$ is the $i_{th}$ element in $W$. For $n$-bit quantization, $Q_{n}=-2^{n-1}$ and $Q_{p}=2^{n-1}-1$ are then determined to constrain the range of quantization, and the scaling factor $s$ can be defined as:
$s=\frac{Q_{p}}{\mathrm{AbsMean}(W)}.$ (3)
Finally, the quantized weights $\tilde{W}$ can be expressed in the following equation:
$\tilde{W}=\mathrm{Clip}[\mathrm{Round}(W\cdot s),Q_{n},Q_{p}]/s,$ (4)
where $\mathrm{Clip}()$ function assures that all the values are in the range of $[Q_{n},Q_{p}]$ and $\mathrm{Round}()$ returns the nearest integer. This allows us to constrain the quantized $\tilde{W}$ to $n$-bits. While for the inputs and activations, we follow the settings introduced in BitNet (Wang et al., 2023; Ma et al., 2024) and quantize them into 8 bits.
After computing the language modeling cross-entropy loss based on $\tilde{W}$ and inputs in each training step, we first allow the optimizer to calculate $W^{\prime}$, the dense weight matrix intended for updating. In traditional QAT, the straight-through estimator is applied, and $W^{\prime}$ replaces the original high-precision $W$, subsequently undergoing the quantization process from Equation (2) to Equation (4) again in the next training step. However, in our approach, we simplify this process by directly applying stochastic rounding on $W^{\prime}$ using Equation (1):
$\widetilde{W}=\mathrm{SR}(W^{\prime}),$ (5)
to make sure it maintains $n$-bits without requiring the retention of high-precision weights, thus getting rid of the straight-through estimator and skipping the process from Equation (2) to Equation (4). During our modified DQT, we directly replace $\tilde{W}$ with $\widetilde{W}$ and proceed to the next training step, ensuring that the weight matrices are always constrained to $n$-bits throughout training. This makes the biggest difference from traditional QAT methods.
4 Experiments
In this section, we begin by detailing our training dataset and implementation setup. Next, we compare models trained with DQT to FP32 models in Section 4.4. We move on to discuss the impact of altering the bit width involved in DQT in Section 4.5. Finally, in Section 4.6, we conduct a fair comparison between our DQT models and BitNet b1.58, both utilizing ternary weight matrices.
4.1 Training Dataset
We use the English Wikipedia dataset (20231101.en) [^2] https://huggingface.co/datasets/wikimedia/wikipedia to pre-train baselines and our models. The maximum length of training data is set to 512. For texts longer than 512, we split them into individual texts while for those shorter than 512, padding is applied. As a result, we acquire a dataset with 14 million sentences from the original 6.4 million examples.
4.2 Baselines
We choose two types of baselines for comparing with our DQT. The first is a reproduced full-precision (FP32) model. We compare DQT to it to demonstrate the effectiveness of our approach with reduced bits. The second baseline is the reproduced BitNet b1.58, a well-known QAT method that employs ternary weights. To ensure a fair comparison with BitNet, we constrain our DQT models to use only ternary weights during inference.
4.3 Implementation Details
We train all our models using fake (simulated) quantization on NVIDIA A100 GPUs to remain hardware-agnostic, without leveraging true low-precision computation. We follow the structure of a 12-layer LLaMA Touvron et al. (2023); Dubey et al. (2024) and initiate our model with the hidden size and the intermediate size set to 768, 2048, respectively. The number of attention heads is set to 12 and the maximum position embeddings follow the maximum input length, which is set to 512. Thus, the total parameters for our model and baselines equal to around 130 million. For the baselines, AdamW is utilized as the optimizer while for DQT models, modified AdamW with stochastic rounding is employed. The cosine scheduler is applied with a 2000 step warm-up. We pre-train all models for one epoch under our dataset on 4 A100 80G GPUs. The batch size is set to 64 on each GPU with no gradient accumulation. The learning rate is grid-searched from {1e-5, 1e-4, 5e-4, 1e-3} for each model. As for the tokenizer, we directly utilize a publicly released pre-trained one  [^3] https://huggingface.co/1bitLLM/bitnet_b1_58-large and do not update it during the training process. We train all of our models under FP32 with mixed precision enabled.
4.4 Training Comparison
Figure 2: Comparison of our DQT and other baselines. The horizontal axis represents the training steps. Optimal learning rates are individually selected for each model, specifically 1e-4 for BitNet and the FP32 model, 1e-3 for DQT 1.58-bit, and 5e-4 for DQT 8-bit implementations.
We present the training loss for DQT variants and our reproduced LLaMA, BitNet b1.58 in Figure 2, using the optimal learning rate for each model. In Figure 2, the blue line indicates the ternary implementation of DQT, demonstrating that it is capable of convergence.
When comparing our DQT ternary variant to the reproduced FP32 model, a performance gap is observed. This is expected, as the DQT model operates with only 1.58-bit information during training, limiting its capacity compared to the FP32 model. However, when longer bit-width models are employed, quantizing weight parameters into 8 bits and maintaining only these 8-bit values during training, the performance becomes much more comparable to the FP32 model.
As for the comparison with BitNet b1.58, since our DQT ternary variant does not rely on high-precision weights during training, it results in lower performance. However, as only ternary weights are maintained throughout the process, this approach potentially requires less memory than BitNet. Even when 8-bit quantization is enabled in DQT, the high-precision information required for BitNet exceeds that of our DQT approach.
4.5 Impact of the Bit Width in DQT
Figure 3: Comparison of different bit widths in DQT. Higher $n$-bit results in better performance. Optimal learning rates are selected for each model: 1e-3 for 1.58-bit and 3-bit models, and 5e-4 for 4-bit and 8-bit models.
Next, we proceed to examine the impact of varying the number of bits in DQT. Specifically, we experiment with $n$ in {1.58, 3, 4, 8}, and the corresponding results are presented in Figure 3. We select the optimal learning rate for each model.
From Figure 3, we can first observe a clear trend: as the number of bits used in DQT increases, the model’s performance improves consistently. Notably, for our 1.58-bit and 3-bit DQT implementation, we can observe some outliers in the blue line and red line, indicating the difficulty of training low-bit models. In contrast, the higher-bit implementations exhibit more stability throughout the training process.
4.6 Inference with Ternary Weights
Since the straight-through estimator is not employed in our proposed DQT, both the forward and backward operates directly on $n$-bits weight matrices , which means that our DQT models trained with larger bit-width are not inherently ternary during inference. To enable a fair comparison with BitNet b1.58, we adapt our models to perform ternary inference. This is achieved by using ternary weights during the forward pass while maintaining $n$-bit weights for the backward pass, updated through the straight-through estimator.
To illustrate the inference performance differences between BitNet b1.58 and our DQT models, we present their loss and perplexity scores by evaluating on the test set of WikiText-2 Merity et al. (2016)  [^4] https://huggingface.co/datasets/Salesforce/wikitext/viewer/wikitext-2-v1/test. The results are shown in Table 1. For consistency, we set the sequence length of WikiText-2 to 512, aligning with our training set. As shown in Table 1, the evaluation loss difference on WikiText-2 between our DQT 8-bit model and BitNet is minimal, with a degradation of about only 5%. When ternary inference is applied, the performance slightly decreases compared to 8-bit inference but remains close, demonstrating the robustness of our approach with ternary inference and its flexibility in deployment.
Methods Loss PPL
FP32 (reproduced) 5.39 41.93
BitNet b1.58 (reproduced) 5.52 45.83
DQT 1.58 bits (ours) 6.20 73.41
DQT 8 bits (ours) 5.80 55.75
DQT 8 bits (ternary Inf.) 5.93 60.98
Table 1: Evaluation results on WikiText-2. Loss and perplexity metrics are reported for comparison. indicates models trained with 8 bits but inference with 1.58 bits.
5 Conclusion
In this work, we explore Direct Quantized Training, a modified QAT method that directly updates low-bit weight matrices without relying on high-precision weights during training, thereby saving memory footprints. Experimental results demonstrate that DQT enables training without updating on high-precision weights, which are required for straight-through estimation. Furthermore, when extending the bit width to 8 bits, DQT models achieve performance levels comparable to FP32 models and BitNet b1.58 in both training and inference.
6 Future Work
First, we plan to extend our experiments with DQT to larger LLMs and datasets as more computational resources become available. We are particularly interested in evaluating the applicability of DQT on larger models, such as 7B parameters.
Next, the current implementation of DQT is simulated on general GPUs. For instance, we apply stochastic rounding on weight matrices optimized by the AdamW optimizer, which retains two high-precision values as momentum and variance for each element in the weight matrices. Besides, the low-bit weight matrices in our experiments are in the FP32 format, following publicly available code implementations [^5] https://huggingface.co/1bitLLM/bitnet_b1_58-large/tree/main to remain GPU-agnostic. As a result, we have not yet observed actual memory savings when comparing DQT to traditional methods like BitNet. Nevertheless, we can confirm that the weight matrices after backpropagation are indeed constrained to $n$-bits, as described in the preceding sections. Furthermore, the stored model parameters are also restricted to $n$-bits, aligning with the design of DQT. Moving forward, we aim to refine the implementation of DQT to achieve genuine memory savings and utilize simpler optimizers, e.g., SGD, to reduce the memory footprints required during optimization.
Besides, while general hardware is mainly optimized for higher-bit arithmetic operations, the support for lower bit-width operations remains sparse. This limits the practical benefits of DQT on existing hardware. In the current implementation of DQT, the low-bit computations are simulated on general GPUs that lack specialized support for ternary or other low-bit arithmetic. If GPUs were equipped with dedicated low-bit arithmetic units designed specifically for ternary or binary operations, we could expect substantial gains in both memory efficiency and computational speed. We welcome further discussion and insights from the community.